Hi please assist with the following questions,

1. My elasticity is 2.0 and I am contemplating a price increase of 20%. If my sales are 1000 right now, how many sales could I expect to gain or lose?

2. My elasticity is .6. What should I do with my price?

3. My sales went up by 100 when I cut my price from $18 to $16. What is my elasticity?
Answer #1


Question 1

Price elasticity of demand = 2

Percentage increase in price = 20%

When the price increases, quantity demanded decreases.

Percentage change in quantity demanded = Price elasticity of demand * Percentage change in price

Percentage change in quantity demanded = 2 * 20 = 40

Thus,

The quantity demanded will decrease by 40%.

Current sales = 1,000 units

Percentage decrease in sales expected = 40%

Decrease in sales = 1,000 * 0.40 = 400 units

Thus,

It is expected that firm will lose sales of 400 units.